# Personalized Speech Emotion Recognition in Human-Robot Interaction using Vision Transformers

### v2

@@ I Introduction
Figure 1: The experimental workflow showing our speech emotion recognition pipeline during human-robot interaction.
Recent advancements in computer vision, particularly with the emergence of Vision Transformers (ViTs), have opened up new possibilities for leveraging visual data in SER [18]. ViTs, originally designed for image classification tasks, have demonstrated exceptional performance in capturing spatial dependencies and global contextual information [19, 20]. Their ability to model long-range dependencies and extract high-level features makes them well-suited for analyzing visual representations of speech, such as spectrograms [21]. The application of ViTs in SER is still an emerging field, and several research directions warrant further exploration.
@@ II Related Works
The authors in [22] proposed to use CNN and RNN pipelines along with data augmentation techniques to improve the robustness of these models. This robustness was crucial for a human-robot interaction scenario with robot’s ego noise. Further, the authors in [23] propose a machine learning pipeline for SER. Their approach involves using personalized and non-personalized features for SER. However, neither of these papers contributes to evaluating transformer-based architectures, which are currently SOTA in numerous fields of study [21].
@@ III-A Data Acquisition
Six neurotypical adults were recruited to participate in a human-robot interaction study to classify their speech into four primary emotions. This one-to-one HRI involves two-way communication between the robot and the participants. Each participant asks pre-defined questions as shown in Figure 1. These questions had been used for our previous studies during HRI [33]. The following are the questions we asked the participants to ask the robot:
@@
The robot responds with appropriate answers to those questions and asks those questions back to the participant. The participants’ replies are not pre-defined. They were asked to reply to the robot’s questions with short answers. For each of these question-and-answer pairs, each participant was asked to speak in an emotional tone depicting one of the four primary emotions, i.e., neutral, happy, sad, and angry. The voices of the participants were recorded during this pseudo-natural human-robot interaction where the questions that the participant asks were pre-defined but their answers weren’t.
@@ III-B Mel Spectrogram
$\begin{split}m=2595\log_{10}\left(1+\frac{f}{700}\right)\\ &=1125\ln\left(1+\frac{f}{700}\right)\end{split}$ (1)
As can be seen form equation 1, the mel scale is a logarithmic scale to convert the frequency of the sounds from Hz to mels. The audio signal first goes through a fast Fourier transform performed on overlapping signal segments. These frequencies are converted to the log scale and the amplitude is converted to decibels to make the color dimension as shown in Figure 1.
@@ III-D Problem Formulation and Proposed Pipeline
The formulation of both of these pipelines remains the same with the only difference of using a pre-trained base ViT encoder (vit-base-patch16-224) for the first pipeline (see Figure 2a) whereas using a base BEiT encoder (microsoft/beit-base-patch16-224-pt22k-ft22k) for pipeline 2 (shown in Figure 2b) [19, 20]. Each image $\bm{x}_{i}^{d}$ is first divided into patches, $x_{p}\in\mathbf{R}^{P\times P\times C}$, where P = 16 is the dimension of the image patch. So the output of the linear projection layer, $x^{\prime}\in\mathbf{R}^{N\times(P^{2}C)}$, where $N$ is the number of patches. The patch and position embedding is then done using:
@@ IV Results and Discussion
We evaluate both the ViT (Figure 2a) and the BEiT (Figure 2b) pipelines in two ways:
@@
We perform full fine-tuning of our models on two A5000 GPUs, using K-Fold-Cross validation (f-fold-cross-validation in our case) with a constant learning rate of $052.00e-$ 05. Further, we evaluate the performance of both pipelines for both Approach 1 and 2 using accuracy, precision, recall, and f-1 scores.
@@
Table II shows the performance of both Approach 1 and Approach 2 for both the ViT and BEiT models. Fine-tuning vanilla ViTs and vanilla BEiTs gives us some promising results when compared to the SOTA. For the RAVDESS dataset, we currently achieve SOTA using the vanilla-ViT model, with the highest performance of 97.49% accuracy as compared to the current SOTA, which has a classification accuracy of 86.70% using multimodal data [25]. For the TESS dataset, we again achieve SOTA using vanilla-ViTs and vanilla-BEiTs, which is very similar to the ones obtained by the authors in [29]. The classification accuracy for the CREMA-D dataset was the highest for the mixed dataset approach (Approach 2) with vanilla-ViTs, which is better than the performance of comparable transformer architectures presented by the authors in [39] and other non-transformer-based approaches [40, 41]. For the ESD dataset, our peak classification accuracy (96.25%) was obtained by a vanilla-BEiT model fine-tuned only on $(\mathcal{X}_{ESD,train},\mathcal{Y}_{ESD,train})$, which is again comparable to the current SOTA (93.20%) as presented by the authors in [32]. Since MELD dataset has numerous speakers, it covers a wide-range of speaker characteristics (see Figure 3). This can be see in the low classification accuracy of the MELD dataset from the Table II. We obtained peak accuracy when the BEiT model fine-tuned over $(\mathcal{X}_{train,mix},\mathcal{Y}_{train,mix})$. However, our results with the MELD come close to the classification accuracies presented by the authors in [26].
TABLE II: Performance of ViT and BEiT pipelines for approach 1 and approach 2 based on accuracy, precision, recall, and F-1 scores.
Dataset Model name 5 fold cross validation learning rate AccuracyPrecision Recall F1-score
RAVDESS google/vit-base-patch16-224 2.00e-05 97.49 0.9749 0.9749 0.9749
@@
microsoft/beit-base-patch16-224-pt22k-ft22k 94.62 0.9486 0.9462 0.9463
TESS google/vit-base-patch16-224 2.00e-05 100.00 1.0 1.0 1.0
microsoft/beit-base-patch16-224-pt22k-ft22k 100.00 1.0 1.0 1.0
CREMA-D google/vit-base-patch16-224 2.00e-05 72.06 0.7237 0.7206 0.7213
microsoft/beit-base-patch16-224-pt22k-ft22k 71.85 0.72 0.7185 0.7173
ESD google/vit-base-patch16-224 2.00e-05 95.84 0.9585 0.9584 0.9584
microsoft/beit-base-patch16-224-pt22k-ft22k 96.25 0.9626 0.9625 0.9625
MELD google/vit-base-patch16-224 2.00e-05 49.83 0.4402 0.4983 0.4601
microsoft/beit-base-patch16-224-pt22k-ft22k 43.32 0.4304 0.4332 0.4317
RAVDESSmix google/vit-base-patch16-224 2.00e-05 95.7 0.9572 0.957 0.957
TESSmix 100.00 1.00 1.00 1.00
CREMA-Dmix 74.51 0.7522 0.7451 0.7467
ESDmix 95.13 0.9513 0.9513 0.9513
MELDmix 49.48 0.4413 0.4948 0.4594
RAVDESSmix microsoft/beit-base-patch16-224-pt22k-ft22k 2.00e-05 94.98 0.9498 0.9498 0.9497
TESSmix 100.00 1.00 1.00 1.00
CREMA-Dmix 72.36 0.7281 0.7236 0.7217
ESDmix 95.28 0.9533 0.9528 0.9528
MELDmix 50.22 0.448 0.5022 0.4638
@@
Figure 3: T-SNE plots of ViT and BEiT embeddings for each emotion of all datasets and our collected participants’ data.
(a) Participant 1
(b) Participant 2
(c) Participant 3
(d) Participant 4
(e) Participant 5
(f) Participant 6
Figure 4: Confusion matrices for participants’ data classification using their best individual models.
@@ IV-A Human subjects’ study
TABLE IV: Model Performance for Six Participants
Participant Name of Model Acc.Pr. Rec. F1 Score Participant Name of Model Acc.Pr. Rec. F1 Score
Participant 1 Vanilla-ViT 56.25 0.5625 0.5625 0.5486 Participant 2 Vanilla-ViT 37.5 0.2905 0.375 0.3189
Vanilla-BEiT 62.5 0..6458 0.625 0.6208 Vanilla-BEiT 43.75 0.333 0.4375 0.3631
ViTmix 68.75 0.7125 0.6875 0.6935 ViTmix 56.25 0.5833 0.5625 0.5607
BEiTmix 75.00 0.8 0.75 0.75 BEiTmix 37.5 0.3854 0.375 0.3512
ViTensemble 68.75 0.7738 0.6875 0.6959 ViTensemble 37.5 0.2708 0.375 0.3006
BEiTensemble 56.25 0.5089 0.5625 0.5284 BEiTensemble 43.75 0.25 0.4375 0.3154
ViTensemble,d 68.75 0.7946 0.6875 0.79 ViTensemble,d 43.75 0.35 0.4375 0.3611
BEiTensemble,d 75.00 0.8304 0.75 0.7193 BEiTensemble,d 56.25 0.6625 0.5625 0.5486
Participant Name of Model Acc.Pr. Rec. F1 Score Participant Name of Model Acc.Pr. Rec. F1 Score
Participant 3 Vanilla-ViT 43.75 0.3571 0.4375 0.3697 Participant 4 Vanilla-ViT 50.00 0.375 0.50 0.4278
Vanilla-BEiT 43.75 0.5769 0.4375 0.3843 Vanilla-BEiT 50.00 0.583 0.5 0.4393
ViTmix 25.00 0.3125 0.25 0.1938 ViTmix 50.00 0.4583 0.5 0.4679
BEiTmix 56.00 0.7875 0.5625 0.5304 BEiTmix 37.5 0.3854 0.375 0.3512
ViTensemble 50.00 0.6111 0.5 0.4622 ViTensemble 75.00 0.7875 0.75 0.7431
BEiTensemble 43.75 0.4405 0.4375 0.4161 BEiTensemble 62.5 0.7986 0.625 0.608
ViTensemble,d 50.00 0.6417 0.5 0.469 ViTensemble,d 56.25 0.525 0.5625 0.5214
BEiTensemble,d 43.75 0.475 0.4375 0.3679 BEiTensemble,d 62.5 0.6667 0.625 0.6071
Participant Name of Model Acc.Pr. Rec. F1 Score Participant Name of Model Acc.Pr. Rec. F1 Score
Participant 5 Vanilla-ViT 37.5 0.211 0.375 0.265 Participant 6 Vanilla-ViT 56.25 0.6 0.5625 0.5754
Vanilla-BEiT 25.00 0.3527 0.25 0.23236 Vanilla-BEiT 43.75 0.4146 0.4375 0.4206
ViTmix 37.5 0.3708 0.375 0.37 ViTmix 37.50 0.4015 0.375 0.3381
BEiTmix 37.5 0.425 0.375 0.3631 BEiTmix 43.75 0.3917 0.4375 0.3944
ViTensemble 43.75 0.333 0.4375 0.375 ViTensemble 62.50 0.70 0.625 0.5972
BEiTensemble 50.00 0.5833 0.5 0.4631 BEiTensemble 50.00 0.5 0.5 0.4667
ViTensemble,d 31.25 0.375 0.3125 0.3167 ViTensemble,d 56.25 0.75 0.5625 0.5916
BEiTensemble,d 43.75 0.4571 0.4375 0.4141 BEiTensemble,d 43.75 0.4208 0.4375 0.4256
@@
We evaluated our speech emotion recognition in a pseudo-naturalistic human-robot interaction scenario using our fine-tuned ViTs and BEiTs. Since each participant asked five questions to the robot and responded to those five questions asked by the robot, we have 40 audio clips from each participant. We divided them into train and test datasets such that two sets of questions and answers each participant gave were separated for the test set. So, each participant had three questions and answer sets for train data. Each of those questions and answers was spoken in a way that depicts each of the four primary emotions of the individual. The split of the train and test data for each participant is shown in Table III. Once the audio has been recorded from the participants, we convert the WAV files into spectrograms as shown in Figure 1.
@@
  • Model 1- Vanilla-ViT and BEiT: Each individual’s data is converted to mel-spectrograms, and then vanilla-ViT and BEiT models are fine-tuned.
  • Model 2- ViTmix and BEiTmix: The fine-tuned models from Approach 2 are fine-tuned on the participants’ mel-spectrograms.
  • Model 3- ViTensemble and BEiTensemble: We use five vanilla-ViTs and five vanilla-BEiTs and average the logits. If the output of each ViTi, where $i=\left\{1,2,3,4,5\right\}$ and of each BEiTi are $c_{i,vit}$ and $c_{i,BEiT}$ respectively, then the ensemble of models is:
@@
  • Model 4- ViTensemble,d and BEiTensemble,d: In this approach, we use the ViTd and BEiTd models trained in Approach 1 on each of the benchmark datasets. So the ensemble works as follows:
@@
Table IV shows the model performance of all the above-proposed models. It becomes evident that the best performance is obtained when we use two broad approaches. First, when we use ViTmix or BEiTmix. This means that fine-tuning a vision transformer-based model on already existing benchmark datasets like the RAVDESS, TESS, CREMA-D, ESD, and MELD helps in recognizing the individual speech emotion characteristics of an individual speaker. The reason can be seen in Figure 3a showing the t-sne plots of the vanilla-ViT embeddings. These t-sne plots show the ViT embeddings for each emotion, plotted for all the benchmark datasets and the participants’ data. The overlap between the speech characteristics of speakers from these benchmark datasets and the participants for our human-robot interaction study helped better classify speech emotion compared to vanilla ViTs or vanilla-BEiTs. For some participants, the ensemble models worked better since their speech characteristics didn’t exactly overlap with the benchmark datasets used in this paper. These outliers can be seen from the t-sne plots in Figure 3. Fine-tuning Vanilla-ViTs or BEiTs on each participant data for recognizing an individual’s speech emotion had the worst classification accuracies as compared to its other counterparts shown in Table III. The confusion matrix for each of the participants’ best performing models have been shown in Figure 4.

### v3

@@ I Introduction
(a) Full Pipeline
(b) Modeli, where $i\in\left\{1,2,3,4\right\}$. Transformer encoder here represents the ViT/BEiT encoder
Recent advancements in computer vision, particularly with the emergence of Vision Transformers (ViTs), have opened up new possibilities for leveraging visual data in SER [18]. This is evident from their superior performance as compared to other deep learning based approaches as shown in [19].
@@
  • We compare our ViT and BEiT models with OpenAI/Whisper-base and ResNet-50 models.
  • We recruited both native and non-native English speakers to include more diverse demographics for robustness.
@@ II Related Works
The authors in [20] proposed to use CNN and RNN pipelines along with data augmentation techniques to improve the robustness of these models. This robustness was crucial for a human-robot interaction scenario with robot’s ego noise. The authors in [21] also used a CNN plus BiLSTM hybrid model for the SER task using SAVEE and TESS datasets. Further, the authors in [22] proposed a machine learning pipeline for SER. Their approach involves using personalized and non-personalized features for SER. However, neither of these papers contributes to evaluating transformer-based architectures, which are currently SOTA in numerous fields of study [23].
@@
For transformer based SER models, some recent works have made attempts to model personalised features of users like the authors in [32]. Other approaches specific to vision transformers based approached for SER include the work by the authors in [33] where they have used the strengths of Multi-Axis Vision Transformer (MaxViT) and the Improved Multiscale Vision Transformer (MViTv2).
@@ III-A Data Acquisition
Twelve neurotypical participants were recruited to participate in a human-robot interaction study to classify their speech into four primary emotions. Six of these participants were native English speakers. The other six were non-native English speakers. This was done to include more diverse demographics to examine SER using vision transformers based models. Among the participants, five were male and the rest were female. All the participants were either students or staff from the university aged between 18-59 years of age. Each participant asks pre-defined questions as shown in Figure 1a. These questions had been used for our previous studies during HRI [34]. The following are the questions we asked the participants to ask the robot:
@@
The robot responds with appropriate answers to those questions and asks those questions back to the participant. The participants’ replies are not pre-defined. They were asked to reply to the robot’s questions with short answers. For each of these question-and-answer pairs, each participant was asked to speak in an emotional tone depicting one of the four primary emotions, i.e., neutral, happy, sad, and angry. The voices of the participants were recorded during this pseudo-natural human-robot interaction where the questions that the participant asks were pre-defined but their answers weren’t. More information on personalization is shared in Algorithm 1.
@@ III-B Mel Spectrogram
$\begin{split}m=2595\log_{10}\left(1+\frac{f}{700}\right)&=1125\ln\left(1+\frac{f}{700}\right)\end{split}$ (1)
As can be seen form equation 1, the mel scale is a logarithmic scale to convert the frequency of the sounds from Hz to mels. The audio signal first goes through a fast Fourier transform performed on overlapping signal segments. These frequencies are converted to the log scale and the amplitude is converted to decibels to make the color dimension as shown in Figure 1a.
@@ III-D Problem Formulation and Proposed Pipeline
The formulation of both of these pipelines remains the same with the only difference of using a pre-trained base ViT encoder (vit-base-patch16-224) for the first pipeline (ViT encoder as the transformer encoder in Figure 1b) whereas using a base BEiT encoder (microsoft/beit-base-patch16-224-pt22k-ft22k) for pipeline 2 (BEiT as the transformer encoder in Figure 1b)[40, 41]. Each image $\bm{x}_{i}^{d}$ is first divided into patches, $x_{p}\in\mathbf{R}^{P\times P\times C}$, where P = 16 is the dimension of the image patch. So the output of the linear projection layer, $x^{\prime}\in\mathbf{R}^{N\times(P^{2}C)}$, where $N$ is the number of patches. The patch and position embedding is then done using:
@@
Algorithm 1 Personalization Process for Speech Emotion Recognition
[…]
0:  y: Accuracy, Precision, Recall, F1 Score, FLOPs, Average Inference Time
1:  Initialization: Prepare robot for interaction.
2:  for each emotion $e\in E$ do
3:     for each question $q\in Q$ do
4:        Person asks the robot question $q$.
5:        Robot responds to question $q$.
6:        Robot asks the same question $q$ back to the person.
7:        Person gives an open-ended reply to question $q$.
8:     end for
9:  end for
10:  Data Collection: Save all responses as audio files (.wav format).
11:  Data Preprocessing:
12:  Convert audio files into mel-spectrograms.
13:  Perform stratified train-test split on the dataset.
14:  Model Fine-Tuning:
15:  Fine-tune the chosen model using the training dataset.
16:  Model Evaluation:
17:  Calculate metrics: Accuracy, Precision, Recall, F1 Score.
18:  Compute FLOPs for one iteration.
19:  Measure average inference time per sample.
20:  Output: Return y
@@ IV Results and Discussion
We evaluate both the ViT and the BEiT pipelines in two ways:
@@
We perform full fine-tuning of our models on two A5000 GPUs, using K-Fold-Cross validation (5-fold-cross-validation in our case) with a constant learning rate of $052.00e-$ 05. Further, we evaluate the performance of both pipelines for both Approach 1 and 2 using accuracy, precision, recall, and f-1 scores.
@@
Table II compares results of Approach 1 and Approach 2 for ViT (ViTd and ViTmix), BEiT (BEiTd and BEiTmix), OpenAI/Whisper-base (openai/whisper-base and openai/whisper-basemix), ResNeT-50 (ResNeT-50d and ResNeT-50mix) models. For the RAVDESS dataset, we currently achieve SOTA using the vanilla-ViT model, with the highest performance of 97.49% accuracy as compared to the current SOTA, which has a classification accuracy of 86.70% using multimodal data [25]. Vanilla-ViT model also outperforms OpenAI/Whisper-base and ResNet-50 models. For the TESS dataset, we again achieve SOTA using vanilla-ViTs and vanilla-BEiTs, which is very similar to the ones obtained by the authors in [19], openai/whisper-base model, and ResNet-50 model. Among our vision transformer based approaches, the classification accuracy for the CREMA-D dataset was the highest for the mixed dataset approach (Approach 2) with vanilla-ViTs, which is better than the performance of comparable transformer architectures presented by the authors in [42] and other non-transformer-based approaches [43, 44]. However, among all the approaches, openai/whisper-base performed the best (80.82%) for the CREMA-D dataset when it was fine-tuned on only the CREMA-D training set. For the ESD dataset, our peak classification accuracy (96.25%) was obtained by a vanilla-BEiT model fine-tuned only on $(\mathcal{X}_{ESD,train},\mathcal{Y}_{ESD,train})$, which is again comparable to the current SOTA (93.20%) as presented by the authors in [31]. It also outperforms openai/whisper-base and ResNet-50 based approach we examined. Since MELD dataset has numerous speakers, it covers a wide-range of speaker characteristics (see Figure 2). This can be see in the low classification accuracy of the MELD dataset from the Table II. Among our ViT and BEiT models, we obtained peak accuracy when the BEiT model fine-tuned over $(\mathcal{X}_{train,mix},\mathcal{Y}_{train,mix})$. However, our results with the MELD come close to the classification accuracies presented by the authors in [26]. In addition to it, based on our experiments, we observed the highest classification accuracies for MELDmix with openai/whisper-base model.
TABLE II: Performance on five emotion datasets for Approach 1 (Unmixed) and Approach 2 (Mixed). “–” indicates unavailable metrics. M1–M5 are explained below. The learning rate used for all the models was 2.00e-5.
Approach 1 (Unmixed Data)
RAVDESS TESS CREMA-D ESD MELD
ID Acc P R F1 Acc P R F1 Acc P R F1 Acc P R F1 Acc P R F1
M1 97.49 0.9749 0.9749 0.9749 100.0 1.0000 1.0000 1.0000 72.06 0.7237 0.7206 0.7213 95.84 0.9585 0.9584 0.9584 49.83 0.4402 0.4983 0.4601
M2 94.62 0.9486 0.9462 0.9463 100.0 1.0000 1.0000 1.0000 71.85 0.7200 0.7185 0.7173 96.25 0.9626 0.9625 0.9625 43.32 0.4304 0.4332 0.4317
M3 84.40 0.8876 0.7905 0.8059 100.0 1.0000 1.0000 1.0000 80.82 0.8103 0.8036 0.8051 97.14 0.9716 0.9714 0.9715 55.97 0.4168 0.3822 0.3925
M4 65.67 0.7002 0.6567 0.6651 100.0 1.0000 1.0000 1.0000 70.41 0.7026 0.7041 0.6999 90.65 0.9081 0.9065 0.9067 51.12 0.4533 0.5112 0.4747
M5 80.60 53.00
Approach 2 (Mixed Data)
RAVDESS_mix TESS_mix CREMA-D_mix ESD_mix MELD_mix
ID Acc P R F1 Acc P R F1 Acc P R F1 Acc P R F1 Acc P R F1
M1 95.70 0.9572 0.9570 0.9570 100.0 1.0000 1.0000 1.0000 74.51 0.7522 0.7451 0.7467 95.13 0.9513 0.9513 0.9513 49.48 0.4413 0.4948 0.4594
M2 94.98 0.9498 0.9498 0.9497 100.0 1.0000 1.0000 1.0000 72.36 0.7281 0.7236 0.7217 95.28 0.9533 0.9528 0.9528 50.22 0.4480 0.5022 0.4638
M3 72.59 0.7873 0.7152 0.7083 100.0 1.0000 1.0000 1.0000 80.00 0.8068 0.7994 0.7993 96.07 0.9612 0.9607 0.9606 56.70 0.4458 0.4021 0.4162
M4 74.63 0.7517 0.7463 0.7474 99.34 0.9936 0.9934 0.9934 74.49 0.7425 0.7449 0.7406 89.44 0.8959 0.8944 0.8946 50.33 0.4747 0.5033 0.4861
M5
Model References:
@@
M1: ViT (google/vit-base-patch16-224)  M2: BEiT (microsoft/beit-base-patch16-224)  M3: Whisper (openAI/whisper-base)  M4: ResNet50  M5: Vesper
@@
Figure 2: T-SNE plots of ViT and BEiT embeddings for each emotion of all datasets and our collected participants’ data. The feature space of the emotional representations for the ViT and the BEiT models for each emotion is shown for all benchmark datasets as well as the participant data.
@@
Remark 1: For a conversational dataset like MELD, methods like meta-learning for few-shot learning, and parameter efficient fine-tuning (PEFT) methods can help learn natural emotions in speech in addition to acted ones [45, 46] for better domain adaptation.
@@ IV-A Human subjects’ study
TABLE IV: Performance Metrics for Each Participant and Model.
Model Mapping: Model 1 - Vanilla-ViT, Model 2 - Vanilla-BEiT, Model 3 - ViTmix, Model 4 - BEiTmix, Model 5 - ViTensemble, Model 6 - BEiTensemble, Model 7 - ViTensemble,d, Model 8 - BEiTensemble,d, Model 9 - openai/Whisper-base, Model 10 - openai/whisper-basemix, Model 11 - resnet-50, Model 12 - resnet-50mix.
Participant Metric Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 Model 10 Model 11 Model 12
1 Accuracy (%) 56.25 62.5 68.75 75.00 68.75 56.25 68.75 75.00 68.75 75.00 50.00 68.75
1 Precision 0.5625 0.6458 0.7125 0.8 0.7738 0.5089 0.7946 0.8304 0.725 0.8166 0.40 0.725
Recall 0.5625 0.625 0.6875 0.75 0.6875 0.5625 0.6875 0.75 0.6875 0.75 0.5 0.6875
F1 Score 0.5486 0.6208 0.6935 0.75 0.6959 0.5284 0.79 0.7193 0.700 0.7166 0.4166 0.70
2 Accuracy (%) 37.5 43.75 56.25 37.5 37.5 43.75 43.75 56.25 43.75 43.75 6.25 25.00
2 Precision 0.2905 0.333 0.5833 0.3854 0.2708 0.25 0.35 0.6625 0.45 0.677 0.0416 0.175
Recall 0.375 0.4375 0.5625 0.375 0.375 0.4375 0.4375 0.5625 0.4375 0.4375 0.0625 0.25
F1 Score 0.3189 0.3631 0.5607 0.3512 0.3006 0.3154 0.3611 0.5486 0.4365 0.425 0.05 0.2055
3 Accuracy (%) 43.75 43.75 25.00 56.00 50.00 43.75 50.00 43.75 25.00 18.75 25.00 31.25
3 Precision 0.3571 0.5769 0.3125 0.7875 0.6111 0.4405 0.6417 0.475 0.196 0.125 0.22 0.2499
Recall 0.4375 0.4375 0.25 0.5625 0.5 0.4375 0.5 0.4375 0.25 0.1875 0.25 0.3125
F1 Score 0.3697 0.3843 0.1938 0.5304 0.4622 0.4161 0.469 0.3679 0.2123 0.1468 0.2197 0.275
4 Accuracy (%) 50.00 50.00 50.00 37.5 75.00 62.5 56.25 62.5 37.5 50.00 25.00 50.00
4 Precision 0.375 0.583 0.4583 0.3854 0.7875 0.7986 0.525 0.6667 .2986 0.5821 0.076 0.4875
Recall 0.5 0.5 0.5 0.375 0.75 0.625 0.5625 0.625 0.375 0.5 0.25 0.5
F1 Score 0.4278 0.4393 0.4679 0.3512 0.7431 0.608 0.5214 0.6071 0.2996 0.4974 0.1176 0.479
5 Accuracy (%) 37.5 25.00 37.5 37.5 43.75 50.00 31.25 43.75 37.5 25.00 25.00 37.5
5 Precision 0.211 0.3527 0.3708 0.425 0.333 0.5833 0.375 0.4571 0.6375 0.1916 0.0625 0.333
Recall 0.375 0.25 0.375 0.375 0.4375 0.5 0.3125 0.4375 0.375 0.25 0.25 0.375
F1 Score 0.265 0.23236 0.37 0.3631 0.375 0.4631 0.3167 0.4141 0.3696 0.2166 0.1 0.29166
6 Accuracy (%) 56.25 43.75 37.50 43.75 62.50 50.00 56.25 43.75 50.00 50.00 31.25 43.75
6 Precision 0.6 0.4146 0.4015 0.3917 0.70 0.5 0.75 0.4208 0.5833 0.6071 0.1905 0.4687
Recall 0.5625 0.4375 0.375 0.4375 0.625 0.5 0.5625 0.4375 0.500 0.500 0.3125 0.4375
F1 Score 0.5754 0.4206 0.3381 0.3944 0.5972 0.4667 0.5916 0.4256 0.4714 0.4864 0.2364 0.4226
7 Accuracy (%) 46.67 33.33 53.33 13.33 46.67 40.00 53.33 46.67 18.75 18.75 31.25 43.75
7 Precision 0.44 0.422 0.4778 0.0667 0.5889 0.544 0.6267 0.4667 0.1548 0.1458 0.2019 0.4688
Recall 0.4667 0.33 0.5333 0.133 0.4667 0.400 0.533 0.4667 0.1875 0.1875 0.3125 0.4375
F1 Score 0.4487 0.3022 0.4857 0.0889 0.4610 0.3859 0.511 0.4222 0.1623 0.1625 0.201 0.4167
8 Accuracy (%) 37.50 25.00 37.50 37.50 37.50 37.50 56.25 31.25 25.00 6.25 25.00 18.75
8 Precision 0.4196 0.1500 0.3750 0.3155 0.4196 0.200 0.5792 0.2458 0.0769 0.0321 0.0625 0.206
Recall 0.3750 0.250 0.3750 0.3750 0.3750 0.3750 0.5625 0.3125 0.250 0.0625 0.2500 0.1875
F1 Score 0.3197 0.1825 0.3416 0.3292 0.3197 0.2540 0.565 0.2736 0.1176 0.0417 0.10 0.1806
9 Accuracy (%) 43.75 56.25 25.00 31.25 43.75 37.50 31.25 37.50 31.25 37.50 18.75 37.50
9 Precision 0.3792 0.4304 0.2167 0.1562 0.4107 0.5089 0.3125 0.3875 0.3083 0.3214 0.1056 0.4167
Recall 0.4375 0.5625 0.2500 0.3125 0.4375 0.3750 0.3125 0.3750 0.3125 0.3750 0.1875 0.3750
F1 Score 0.4042 0.4804 0.2306 0.2083 0.4205 0.3784 0.3054 0.3681 0.2944 0.3409 0.1325 0.3667
10 Accuracy (%) 37.50 37.50 56.25 43.75 50.00 62.50 37.50 56.25 18.75 56.25 25.00 31.25
10 Precision 0.300 0.2798 0.5458 0.4405 0.500 0.6875 0.3583 0.5833 0.1458 0.4446 0.0667 0.400
Recall 0.3750 0.3750 0.5625 0.4375 0.500 0.6250 0.3750 0.5625 0.1875 0.5625 0.25 0,3125
F1 Score 0.3056 0.3123 0.5506 0.4161 0.4921 0.6446 0.3631 0.5357 0.1548 0.4905 0.1053 0.333
11 Accuracy (%) 60.00 46.67 60.00 73.33 53.33 40.00 60.00 66.67 60.00 73.33 40.00 40.00
11 Precision 0.6457 .2519 0.7852 0.7467 0.5733 0.3556 0.5778 0.7968 0.6250 0.7875 0.2067 0.5378
Recall 0.600 0.4667 0.600 0.7333 0.533 0.400 0.6 0.6667 0.5833 0.7292 0.400 0.400
F1 Score 0.5606 0.3241 0.5708 0.7304 0.4590 0.3111 0.5511 0.6139 0.5893 0.7411 0.2667 0.4394
12 Accuracy (%) 37.50 43.75 50.00 31.25 62.50 50.00 62.50 43.75 31.25 37.50 31.25 37.50
12 Precision 0.433 0.6458 0.6071 0.2979 0.7778 0.4970 0.6417 0.3611 0.3006 0.3083 0.3167 0.4437
Recall 0.3750 0.4375 0.500 0.3125 0.6250 0.500 0.6250 0.4375 0.3125 0.3750 0.3125 0.3750
F1 Score 0.3265 0.4446 0.4697 0.2956 0.6300 0.4705 0.6290 0.3681 0.2963 0.3361 0.2053 0.3361
@@
We evaluated our speech emotion recognition in a pseudo-naturalistic human-robot interaction scenario using our fine-tuned ViTs and BEiTs. Since each participant asked five questions to the robot and responded to those five questions asked by the robot, we have 40 audio clips from each participant. We divided them into train and test datasets such that two sets of questions and answers each participant gave were separated for the test set. So, each participant had three questions and answer sets for train data. Each of those questions and answers was spoken in a way that depicts each of the four primary emotions of the individual. The split of the train and test data for each participant is shown in Table III. Once the audio has been recorded from the participants, we convert the WAV files into spectrograms as shown in Figure 1a.
@@
  • Model 1 and 2- Vanilla-ViT and BEiT: Each individual’s data is converted to mel-spectrograms, and then vanilla-ViT and BEiT models are fine-tuned.
  • Model 2 and 3- ViTmix and BEiTmix: The fine-tuned models from Approach 2 are fine-tuned on the participants’ mel-spectrograms.
  • Model 3 and 4- ViTensemble and BEiTensemble: We use five vanilla-ViTs and five vanilla-BEiTs and average the logits. If the output of each ViTi, where $i=\left\{1,2,3,4,5\right\}$ and of each BEiTi are $c_{i,vit}$ and $c_{i,BEiT}$ respectively, then the ensemble of models is:
@@
  • Model 5 and 6- ViTensemble,d and BEiTensemble,d: In this approach, we use the ViTd and BEiTd models trained in Approach 1 on each of the benchmark datasets. So the ensemble works as follows:
@@
TABLE V: Model complexity and inference times
Model FLOPs Inference Time (ms/sample)
Vanilla-ViT (Model 1) 16.87 GMac 0.516
Vanilla-BEiT (Model 2) 17.59 GMac 3.315
ViTmix (Model 3) 16.87 GMac 0.4684
BEiTmix (Model 4) 17.59 GMac 3.339
ViTensemble (Model 5) 84.34 GMac 2.418
BEiTensemble (Model 6) 87.94 GMac 16.6026
ViTensemble,d (Model 7) 84.34 GMac 2.188
BEiTensemble,d (Model 8) 87.94 GMac 13.581
OpenAI/Whisper-base (Model 9) 30.11 GMac 197.141
OpenAI/Whisper-basemix (Model 10) 30.11 GMac 197.141
ResNet-50mix (Model 11) 4.13 GMac 0.517
ResNet-50mix (Model 12) 4.13 GMac 0.508
Table IV shows the model performance of all the above proposed models. It becomes evident that the best performance is obtained when we use ViT or BEiT based approaches as compared to OpenAi/Whisper-base and ResNet-50. As can be seen from Figure 2a and 2b, the participant data has an overlap in the feature space of the datasets used in this paper. The overlap between the speech characteristics of speakers from these benchmark datasets and the participants for our human-robot interaction study helped better classify speech emotion compared to vanilla ViTs or vanilla-BEiTs. This contributes to the participants having better classification accuracies for the mix models and the ViTensemble,d/BEiTensemble,d (see Table IV) for participants 1, 2, 3, 7, 8, 11, and 12. For some participants, the ensemble models (ViTensemble and BEiTensemble) worked better since their speech characteristics didn’t exactly overlap with the benchmark datasets used in this paper. For both native and non-native English speakers, ViT and BEiT based models performed better than other models compared. For time complexity and inference times of our models, we analysed Floating Point Operations (FLOPs) and also recorded the average time it takes each of our models to classify one input test sample. As can be seen from Table IV, all of the participants had the best classification accuracies with either a ViT or BEiT based model except for participant 11, who had the same accuracy for the openai/whisper-base model too. However, the inference time for the openai/whisper-base was significantly higher (197.141 ms/sample) than the BEiTmix model (3.339 ms/sample). Note that, real-time deployment of SER systems for HRI also depends on the system-specific requirements.